A rock band plays at a 67 dB sound level. How many times greater is the intensity from another rock band playing at 105 dB?
Answer #1

here,

for dB1 = 67 dB

let the intensity of sound level be dB1

dB = 10 * log(I1 /I0)

67 = 10 * log(I1 /(1 * 10^-12))

solving for I1

I1 = 5.01 * 10^-6 W/m^2

for dB2 = 105 dB

let the intensity of sound level be dB2

dB = 10 * log(I2 /I0)

105 = 10 * log(I2 /(1 * 10^-12))

solving for I2

I2 = 3.16 * 10^-2 W/m^2

I2 /I1 = (3.16 * 10^-2) / (5.01 * 10^-6)

I2 /I1 = 6.31 * 10^3